OpenStudy (anonymous):

How do I solve the integral of (3t-2)/(t+1) dt ?

OpenStudy (anonymous):

do long division..:)

OpenStudy (anonymous):

Try \(u=t+1\)

OpenStudy (anonymous):

\[ 3t-2 = 3t - 3-2+3 = 3(t-1)+1 = 3u+1 \]

OpenStudy (kenljw):

(3t-2)/(t+1) = A + B/(t + 1) A (t + 1) + B = 3t - 2 : A = 3, B = -5 now it's easily integrated

OpenStudy (anonymous):

i thought u = t+1, not t-1

OpenStudy (anonymous):

@Yahoo! has it right... divide then integrate